# Causal Models in Requirement Specifications for Machine Learning: A vision

### v1

@@ Abstract.
copyright: cc
@@ 3. Causal modelling as an RE activity
In a typical ML development pipeline, causal modelling would be a step between problem definition and data collection as it allows to formalise domain knowledge, identify relevant variables, and refine data requirements by distinguishing causal relationships from mere correlations before collecting the training data. Particularly, graphical causal models in the form of directed acyclic graph (DAG) allow to communicate explicitly assumed directions of causality and assumptions about confounders, i.e., situations in which a variable $Z$ is associated to two unrelated random variables $X_{1}$ and $X_{2}$ such that a spurious relationship between $X_{1}$ and $X_{2}$ can be observed: $X_{1}\,\leftarrow\,Z\,\rightarrow\,X_{2}$.
@@
(1) $p(X_{1},\ldots,X_{n})=\prod_{i=1}^{n}p\left(X_{i}|\mathbf{PA}_{i}\right)$
Causal factorisation implies that an observed joint distribution of interest can be decomposed into a product of conditional distributions, where each term corresponds to a causal mechanism.
Step 1) Identify individual causal mechanisms: The idea is to identify individual causal mechanisms based on high-level requirements, prior domain knowledge, and context assumptions.
@@ 4. Demonstration on industrial prototype
We held three workshops with two Siemens engineers and two academic researchers to explore using causal models for requirements specification in the second half of 2022. The researchers introduced causal models with examples like temperature $\leftarrow$ sunrise $\rightarrow$ birds chirping and explaining key concepts such as confounding, colliders, and d-separation using for example the back-door criteria. The company experts then presented the prototype system, and prior knowledge rules were formalised. Together, the team identified causal mechanisms and updated the causal model iteratively with each newly found causal mechanism. We then applied d-separation to close non-causal paths which resulted in data and model requirements to ensure the ML model controls for potential confounding.
@@ Results:
As an example for a resulting test case, ID1 states that classification is independent of $T_{E}$ given $H_{s}$, $T_{s}$, and $V_{s}$. A test case could trigger faults at varying $T_{E}$ to verify that the detection probability remains unchanged. As an example for runtime monitoring, ID5 suggests $V_{s}$ and $T_{E}$ should be independent. A monitor could track their correlation and trigger an alarm if a threshold is exceeded which would indicate a shift in the assumed operational context (e.g., the probability of a mechanical fault could depend on the environmental temperature which would be a violation of the assumed causal models for this system).
@@ 5. Discussion and research agenda
Completeness criteria for variable selection. A key challenge is knowing when a causal graphs includes “enough” prior knowledge. We need methods to identify the necessary set of variables that must be included for a given use case and which completeness criteria can guide the search for a sufficient causal representation.
@@
Testing and runtime checks. ML software system must align with expected (causal) behaviour. Causal graphs imply independence criteria that lead to testable implications for the runtime behaviour. Research should explore how to translate these into testing strategies and monitors and how reliable such monitors are in practise.
@@ Conclusion
Causal reasoning offers a systematic way to integrate prior knowledge into RE for ML software systems. We outlined a vision and demonstrated a preliminary workflow to derive and argue for low level model and data requirements from high level prior knowledge using causal graphs. We discussed future research activities that are needed to turn this vision into industrial practise.

### v2

@@ Abstract.
copyright: ccccs: Software and its engineering Requirements analysisccs: Software and its engineering Reusabilityccs: Software and its engineering Software prototyping
@@ 3. Causal modelling as an RE activity
In a typical ML development pipeline, causal modelling would be a step between problem definition and data collection as it allows to formalise domain knowledge, identify relevant variables, and refine data requirements by distinguishing causal relationships from mere correlations before collecting the training data. Particularly, graphical causal models in the form of directed acyclic graph (DAG) allow to communicate explicitly assumed directions of causality and assumptions about confounders, i.e., situations in which a variable $Z$ is associated to two random variables $X_{1}$ and $X_{2}$ such that a spurious relationship between $X_{1}$ and $X_{2}$ can be observed: $X_{1}\,\leftarrow\,Z\,\rightarrow\,X_{2}$.
@@
(1) $p(X_{1},\ldots,X_{n})=\prod_{i=1}^{n}p\left(X_{i}|\mathbf{PA}_{i}\right),$
where $\mathbf{PA}_{i}$ denotes the set of parents (variables that have a direct causal effect) of a variable $X_{i}$ in the DAG. Causal factorisation implies that an observed joint distribution of interest can be decomposed into a product of conditional distributions, where each term corresponds to a causal mechanism.
Step 1) Identify individual causal mechanisms: The aim is to pinpoint specific cause-effect pathways informed based on high-level requirements, prior domain knowledge, and context assumptions.
@@ 4. Demonstration on industrial prototype
We held three workshops with two Siemens engineers and two academic researchers to explore using causal models for requirements specification in the second half of 2022. The researchers introduced causal models with examples like temperature $\leftarrow$ sunrise $\rightarrow$ birds chirping and explaining key concepts such as confounding, colliders, and d-separation using for example the back-door criteria. The company experts then presented the prototype system, and prior knowledge rules were formalised together by identifying causal mechanisms and updating the causal model iteratively with each newly found causal mechanism. We then applied d-separation to close non-causal paths between the exposure (i.e., a cooling fault) and the outcome (i.e., the classification result) which resulted in data and model requirements to ensure the ML model controls for potential confounding.
@@ Results:
As an example for a resulting test case, ID1 states that classification is independent of $T_{E}$ given $H_{s}$, $T_{s}$, and $V_{s}$. A test case could trigger faults at varying $T_{E}$ to verify that the detection probability remains unchanged. As an example for runtime monitoring, ID5 suggests $V_{s}$ and $T_{E}$ should be independent. An additional monitor could track their correlation during operation and trigger an alarm if a threshold is exceeded which would indicate a shift in the assumed operational context (e.g., the probability of a mechanical fault could depend on the environmental temperature which would be a violation of the assumed causal models for this system).
@@ 5. Discussion and research agenda
Criteria for sufficient variable selection. A key challenge is knowing when a causal graphs includes “enough” prior knowledge. We need methods to determine a sufficient set of variables that must be included for a given use case and methods for deciding between competing causal DAGs given a ´´sufficient” set of variables.
@@
Testing and runtime checks. ML software system must align with expected (causal) behaviour. Causal graphs imply independence criteria that lead to testable implications for the runtime behaviour. Research should explore how to translate these into testing strategies and monitors and how reliable such monitors are in practice.
@@ Conclusion
Causal reasoning offers a systematic way to integrate prior knowledge into RE for ML software systems. We outlined a vision and demonstrated a preliminary workflow to derive and argue for low level model and data requirements from high level prior knowledge using causal graphs. We discussed future research activities that are needed to turn this vision into industrial practice.
Acknowledgements.
This project has received funding from the EU’s Horizon research and innovation program under grant agreement No 957197 (vedliot).